# Bondi mass, memory effect and balance law of polyhomogeneous spacetime

### v2

@@ Abstract.
Spacetimes with metrics admitting an expansion in terms of a combination of powers of $1/r$ and $\ln r$ are known as polyhomogeneous spacetimes. The asymptotic behaviour of the Newman-Penrose quantities for these spacetimes is presented under certain gauges. The Bondi mass is revisited via the Iyer-Wald formalism. The memory effect of the gravitational radiation in the polyhomogeneous spacetimes is also discussed. It is found that the appearance of the logarithmic terms does not affect the balance law and it remains unchanged as the one of spacetimes with metrics admitting an expansion in terms of powers of $1/r$.
Key words and phrases:
Polyhomogeneous spactimes, Newman-Penrose formalism, Bondi mass, memory effect, balance law
2010 Mathematics Subject Classification:
83C60 (Primary); 83C22 (Secondary)
1. Introduction
@@
Nearly a century ago, Einstein predicted gravitational waves based on an analysis of the linearised field equations of general relativity. However, due to the complexity of diffeomorphisms, the existence of gravitational waves caused a huge controversy for quite a long time [1].
@@
Bondi et al. solved this debate theoretically by proposing the Bondi-Sachs (BS) framework [2, 3]. The asymptotic behaviour of metric functions in the BS coordinate system is obtained by solving the Einstein field equations using the formal series expansions in powers of $1/r$, where the coordinate $r$ denotes the luminosity distance. Within the BS framework, the Bondi mass was successfully defined and it satisfied the well-known mass loss formula [2, 3]. The Bondi-Metzner-Sachs (BMS) group is the asymptotic symmetry group and the Bondi mass possesses a uniqueness property [4]. Newman, Penrose and Unti also used the formal $1/r$ expansions with $r$ the affine parameter of a future-oriented null geodesic to study the asymptotically flat spacetimes within the Newman-Penrose (NP) formalism [5, 6, 7, 8]. Bondi et al. excluded the logarithmic terms in the asymptotic expansions by precluding the appearance of $1/r^{2}$ terms in some of the metric functions.This assumption is known as the outgoing radiation condition. Newman et al. used the condition that the asymptotic behaviour of the $\Psi_{0}$ component of the Weyl tensor is $O(r^{-5})$. If a slower decay assumption for $\Psi_{0}$ is adopted, the logarithmic terms will appear in the asymptotic expansions. In this paper, we referred to a spacetime which can be asymptotically expanded in terms of $1/r$ as a smooth asymptotically flat spacetime.
@@
In the celebrated work of the nonlinear stability of the Minkowski spacetime, Christodoulou and Klainerman proved that a generic asymptotically flat spacetime may admit a slower decay with $\Psi_{0}$ falling off as $O(r^{-7/2})$ [9]. This scenario strongly suggests that the purely $1/r$ expansions are not adequately general. The limitations of the $1/r$ expansions have been noticed by many researchers and it seems natural to combine the powers of both $1/r$ and $\ln r$ in the asymptotic expansions [10, 11, 12, 13, 14, 15]. Roughly speaking, a spacetime which can be expanded asymptotically in terms of $r^{-i}(\ln r)^{j}$ is called a polyhomogeneous spacetime, cf. [13] for the precise definition. Chruściel, MacCallum and Singleton systematically analyzed the structure of polyhomogeneous spacetimes by a Bondi-Sachs type method [13]. They showed that the assumption of polyhomogeneity is formally consistent with the Einstein equations. Later, Valiente Kroon managed to construct the logarithmic Newman-Penrose constants for polyhomogeneous spacetimes [14, 15]. Gasperin and Valiente Kroon investigated the relations between the polyhomogeneous expansions and the freely specifiable parts of time symmetric initial data [16]. Fursaev et al. found that a spacetime generated by the cosmic strings admits asymptotically polyhomogeneous form [17]. Recently, Geiller, Laddha and Zwikel studied the symmetries of the gravitational scattering in the absence of peeling within the Bondi-Sachs framework [18].
@@
The notion of mass in general relativity is a quite subtle issue. Iyer and Wald established an effective framework to naturally introduce conserved quantities, including the mass [19]. This formalism works well in various diffeomorphism-covariant theories of gravity. Within the Iyer-Wald formalism, it was shown that the Arnowitt-Deser-Misner (ADM) mass is simply the Hamiltonian conjugate to the asymptotic time translation in spacetimes that are asymptotically flat at spatial infinity. To study the conserved quantities at null infinity, Wald et al. introduced a modified Hamiltonian and discussed the BMS charges in the conformal (unphysical) spacetime [20, 21]. In the current paper, we reexamine the Noether charge associated with the asymptotic time translation at null infinity in a direct way in the physical spacetime and no conformal compactification is required. It is shown that, for smooth asymptotically flat spacetimes, the Noether charge associated with the asymptotic time translation turns out to be the Bondi mass in the BS framework and the Newman-Unti (NU) mass respectively in the Newman-Penrose formalism. Moreover, the notion of mass at null infinity in polyhomogeneous spacetimes is discussed.
@@
Gravitational wave memory is a significant prediction of general relativity which was firstly found by Zel’dovich and Polnarev [22]. This type of memory is produced by the change of quadrupole moments and is referred to as the linear memory. In the 1990s, Christodoulou and Frauendiener found the nonlinear memory of gravitational waves [23, 24]. Recently, for smooth asymptotically flat spacetimes, Nichols et al. proposed a BMS method to calculate the gravitational wave memory [25, 26]. Cao et al. applied the BMS method to calculate the gravitational memory produced by binary black hole coalescence [27, 28]. One of the essential steps in the BMS method is the derivation of the balance law [27, Eq.(15)]. The gravitational wave memory and the balance law for polyhomogeneous spacetimes are discussed in this paper. It is shown that the appearance of the logarithmic terms does not affect the balance law and it remains unchanged as the one that for smooth asymptotically flat spacetimes.
@@
We adopt the geometric units $c=G=1$ throughout this paper. The mostly minus signature is used in this work except for Section 4 where the mostly plus signature is adopted following the convention used by Bondi et al. in [2, 3, 29] .
2. Asymptotic behaviour of Newman-Penrose quantities
@@
We follow the notions in [5]. In addition to the null vector $l^{a}$, another null vector $n^{a}$ and the other two complex null vectors $m^{a},\bar{m}^{a}$ can be written as
@@
$\kappa=\epsilon=\pi=0,\ \rho=\bar{\rho},\ \tau=\bar{\alpha}+\beta.$ (2.3)
This gauge condition represents the fact that $l_{a}=\nabla_{a}u$ is a gradient field and one requires that $m^{a}$ and $n^{a}$ be parallelly propagated along $l^{a}$. The gauge choice (2.3) we adpot here is slightly different with the one used in [15, Eqs.(13) and (14)].
Within the Newman-Penrose formalism, the asymptotic characteristic initial-value problem [5] is usually set as
(i) $\Psi_{0}$ on an initial null hypersurface $\mathcal{N}_{0}$;
(ii) $\sigma$ on $\mathscr{I}^{+}$;
(iii) $\Psi_{1}$, $\rm{Re}\Psi_{2}$ on $\mathscr{I}^{+}\cap\mathcal{N}_{0}.$
The most general polyhomogeneous form for the physically reasonable $\Psi_{0}$ [15] is
@@
The 38 NP equations (including the metric equations, the spin coefficient equations and the Bianchi identities) can be solved order-by-order. When solving the NP equations, one usually needs to calculate both the derivatives and the products of the polyhomogeneous functions [14]. Let $h$ and $g$ be two polyhomogeneous functions as
@@
where $\rho_{i,k}$ and $\sigma_{i,k}$ are functions of $u,\theta$ and $\varphi$. We will denote by ##\## the degree of a polynomial in $z$ and use the convention that the zero polynomial has degree $\infty$- ∞.
In the following, we will solve the NP equations order-by-order. The procedure is similar to the one done in [5, 15]. We only give the details for the solving process of the first two equations in the NP hierarchy which are
@@
$\begin{split}\bullet\ \sigma=&\frac{\sigma_{2}(z)}{r^{2}}+\frac{\sigma_{3}(z)}{r^{3}}+\frac{\sigma_{4}(z)}{r^{4}}+\cdots,\\ &{}^{\#}\sigma_{2}=N_{3}+1,\ ^{\#}\sigma_{3}=N_{4},\ ^{\#}\sigma_{4}=\max\{3N_{3}+3,N_{5}\},\cdots,\\ &\sigma_{2}=\sigma_{2,0}(u,\theta,\varphi)+\sum_{i=1}^{N_{3}+1}\sigma_{2,i}(\theta,\varphi)(\ln r)^{i},\\ &\sigma_{2,j}=\frac{1}{j}\Psi_{0}^{3,j},\ (j=1,2,\cdots,N_{3}+1).\end{split}$ (2.29)
$\begin{split}\bullet\ \rho=&-\frac{1}{r}+\frac{\rho_{3}(z)}{r^{3}}+\frac{\rho_{4}(z)}{r^{4}}+\cdots,\\ &{}^{\#}\rho_{3}=2N_{3}+2,\ ^{\#}\rho_{4}=N_{3}+N_{4}+1,\cdots,\\ &\rho_{3}=\sum_{j=0}^{2N_{3}+2}\rho_{3,j}(\ln r)^{j},\\ &\rho_{3,2N_{3}+2}=-\sigma_{2,N_{3}+1}\bar{\sigma}_{2,N_{3}+1},\\ &\rho_{3,j}=(j+1)\rho_{3,j+1}-\sum_{k=0}^{j}\sigma_{2,k}\bar{\sigma}_{2,j-k},\ (j=2N_{3}+1,\cdots,1,0).\end{split}$ (2.30)
@@
$\begin{split}\bullet\ \alpha=&\frac{\alpha_{1}}{r}+\frac{\alpha_{2}(z)}{r^{2}}+\cdots,\ \ \ ^{\#}\alpha_{1}=0,\ ^{\#}\alpha_{2}=N_{3}+1,\cdots,\\ &\alpha_{1}=\alpha_{1,0}=-\frac{\cot\theta}{2\sqrt{2}},\\ &\alpha_{2}=\sum_{j=0}^{N_{3}+1}\alpha_{2,j}(\ln r)^{j},\\ &\alpha_{2,N_{3}+1}=-\beta_{1,0}\bar{\sigma}_{2,N_{3}+1},\ \ \beta_{1,0}=\frac{\cot\theta}{2\sqrt{2}},\\ &\alpha_{2,j}=(j+1)\alpha_{2,j+1}-\beta_{1,0}\bar{\sigma}_{2,j},\ (j=N_{3},\cdots,1,0).\ \ \ \ \ \ \ \ \ \ \ \ \ \end{split}$ (2.31)
@@
$\begin{split}\bullet\ \beta=&\frac{\beta_{1}}{r}+\frac{\beta_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\beta_{1}=0,\ ^{\#}\beta_{2}=N_{3}+1,\cdots,\\ &\beta_{1}=\beta_{1,0}=\frac{\cot\theta}{2\sqrt{2}},\\ &\beta_{2}=\sum_{j=0}^{N_{3}+1}\beta_{2,j}(\ln r)^{j},\\ &\beta_{2,N_{3}+1}=-\alpha_{1,0}\sigma_{2,N_{3}+1},\\ &\beta_{2,j}=(j+1)\beta_{2,j+1}-\alpha_{1,0}\sigma_{2,j}-\Psi_{1}^{3,j},\ (j=N_{3},\cdots,1,0).\ \ \ \end{split}$ (2.32)
$\begin{split}\bullet\ \tau=&\frac{\tau_{2}}{r^{2}}+\frac{\tau_{3}}{r^{3}}+\cdots,\ \ \ ^{\#}\tau_{2}=N_{3},\cdots,\\ &\tau_{2}=\sum_{j=0}^{N_{3}}\tau_{2,j}(\ln r)^{j},\\ &\tau_{2,j}=\bar{\alpha}_{2,j}+\beta_{2,j},\ (j=0,1,\cdots,N_{3}).\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \end{split}$ (2.33)
$\begin{split}\bullet\ \lambda=&\frac{\lambda_{1}}{r}+\frac{\lambda_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\lambda_{1}=0,\ ^{\#}\lambda_{2}=N_{3}+1,\cdots,\\ &\lambda_{1}=\dot{\bar{\sigma}}_{2,0},\\ &\lambda_{2}=\sum_{j=0}^{N_{3}+1}\lambda_{2,j}(\ln r)^{j},\\ &\lambda_{2,N_{3}+1}=-\mu_{1,0}\bar{\sigma}_{2,N_{3}+1}=\frac{1}{2}\bar{\sigma}_{2,N_{3}+1},\\ &\lambda_{2,j}=(j+1)\lambda_{2,j+1}+\frac{1}{2}\bar{\sigma}_{2,j},\ (j=N_{3},\cdots,1,0).\ \ \ \ \ \ \ \ \ \ \ \ \ \end{split}$ (2.34)
$\begin{split}\bullet\ \mu=&\frac{\mu_{1}}{r}+\frac{\mu_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\mu_{1}=0,\ ^{\#}\mu_{2}=N_{3}+1,\cdots,\\ &\mu_{1}=-\frac{1}{2},\\ &\mu_{2}=\sum_{j=0}^{N_{3}+1}\mu_{2,j}(\ln r)^{j},\\ &\mu_{2,N_{3}+1}=-(\lambda_{1,0}\sigma_{2,N_{3}+1}+\Psi_{2}^{3,N_{3}+1}),\\ &\mu_{2,j}=(j+1)\mu_{2,j+1}-\sigma_{2,j}\lambda_{1,0}-\Psi_{2}^{3,j},\ (j=N_{3},\cdots,1,0).\end{split}$ (2.35)
@@
$\begin{split}\bullet\ \gamma=&\frac{\gamma_{2}}{r^{2}}+\frac{\gamma_{3}}{r^{3}}+\cdots,\ \ \ \ ^{\#}\gamma_{2}=N_{3}+1,\cdots,\\ &\gamma_{2}=\sum_{j=0}^{N_{3}+1}\gamma_{2,j}(\ln r)^{j},\\ &\gamma_{2,N_{3}+1}=-\frac{1}{2}\Psi_{2}^{3,N_{3}+1},\\ &\gamma_{2,j}=\frac{1}{2}\bigg{[}(j+1)\gamma_{2,j+1}-\alpha_{1,0}\tau_{2,j}-\beta_{1,0}\bar{\tau}_{2,j}-\Psi_{2}^{3,j}\bigg{]},(j=N_{3},\cdots,1,0).\end{split}$ (2.36)
$\begin{split}\bullet\ \nu=&\frac{\nu_{1}}{r}+\frac{\nu_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\nu_{1}=0,\ ^{\#}\nu_{2}=N_{3}+1,\cdots,\\ &\nu_{1}=-\Psi_{3}^{2,0}=\eth\dot{\bar{\sigma}}_{2,0},\\ &\nu_{2}=\sum_{j=0}^{N_{3}+1}\nu_{2,j}(\ln r)^{j},\\ &\nu_{2,N_{3}+1}=-\frac{1}{2}\Psi_{3}^{3,N_{3}+1},\\ &\nu_{2,j}=\frac{1}{2}\bigg{[}(j+1)\nu_{2,j+1}+\frac{1}{2}\bar{\tau}_{2,j}-\lambda_{1,0}\tau_{2,j}-\Psi_{3}^{3,j}\bigg{]},\ (j=N_{3},\cdots,1,0).\end{split}$ (2.37)

### v3

@@ Abstract
Spacetimes with metrics admitting an expansion in terms of a combination of powers of $1/r$ and $\ln r$ are known as polyhomogeneous spacetimes. The asymptotic behaviour of the Newman-Penrose quantities for the vacuum polyhomogeneous spacetimes is presented under certain gauges. The Bondi mass is revisited via the Iyer-Wald formalism. The memory effect of the gravitational radiation in the vacuum polyhomogeneous spacetimes is also discussed. It is found that the appearance of the logarithmic terms does not affect the balance law and it remains identical to that of spacetimes with metrics admitting an expansion in terms of powers of $1/r$.
pacs:
04.30.-w, 04.20.-q
1 Introduction
@@
Over a century ago, Einstein predicted gravitational waves based on an analysis of the linearised field equations of general relativity Einstein1916 . However, due to the complexity of diffeomorphisms, the existence of gravitational waves caused a huge controversy for quite a long time Pirani1956 ; Ken07 .
@@
Bondi et al. solved this debate theoretically by proposing the Bondi-Sachs (BS) framework Bondi62 ; Sachs62 . The asymptotic behaviour of metric functions in the BS coordinate system is obtained by solving the Einstein field equations using the formal series expansions in powers of $1/r$, where the coordinate $r$ denotes the luminosity distance. Within the BS framework, the Bondi mass was successfully defined and it satisfied the well-known mass loss formula Bondi62 ; Sachs62 . The Bondi-Metzner-Sachs (BMS) group is the asymptotic symmetry group and the Bondi mass possesses a uniqueness property Chru98 . Newman, Penrose and Unti also used the formal $1/r$ expansions with $r$ the affine parameter of a future-oriented null geodesic to study the asymptotically flat spacetimes within the Newman-Penrose (NP) formalism NP62 ; NU62 ; NP65 ; NP68 . Bondi et al. excluded the logarithmic terms in the asymptotic expansions by precluding the appearance of $1/r^{2}$ terms in some of the metric functions.This assumption is known as the outgoing radiation condition. Newman et al. used the condition that the asymptotic behaviour of the $\Psi_{0}$ component of the Weyl tensor is $O(r^{-5})$. If a slower decay assumption for $\Psi_{0}$ is adopted, the logarithmic terms will appear in the asymptotic expansions. In this paper, we referred to a spacetime which can be asymptotically expanded in terms of $1/r$ as a smooth asymptotically flat spacetime.
@@
In the celebrated work of the nonlinear stability of the Minkowski spacetime, Christodoulou and Klainerman proved that a generic asymptotically flat spacetime may admit a slower decay with $\Psi_{0}$ falling off as $O(r^{-7/2})$ Christo1993 . This scenario strongly suggests that the purely $1/r$ expansions are not adequately general. The limitations of the $1/r$ expansions have been noticed by many researchers and it seems natural to combine the powers of both $1/r$ and $\ln r$ in the asymptotic expansions Fock57 ; Bicak93 ; Ashtekar97 ; Chrusciel95 ; Kroon98 ; Kroon99 . Roughly speaking, a spacetime which can be expanded asymptotically in terms of $r^{-i}(\ln r)^{j}$ is called a polyhomogeneous spacetime, cf. Chrusciel95 for the precise definition. Chruściel, MacCallum and Singleton systematically analyzed the structure of polyhomogeneous spacetimes by a Bondi-Sachs type method Chrusciel95 . They showed that the assumption of polyhomogeneity is formally consistent with the Einstein equations. Later, Valiente Kroon managed to construct the logarithmic Newman-Penrose constants for polyhomogeneous spacetimes Kroon98 ; Kroon99 . Gasperin and Valiente Kroon investigated the relations between the polyhomogeneous expansions and the freely specifiable parts of time symmetric initial data Kroon17 . Fursaev et al. found that a spacetime generated by the cosmic strings admits asymptotically polyhomogeneous form Fursaev23 . Recently, Geiller, Laddha and Zwikel studied the symmetries of the gravitational scattering in the absence of peeling within the Bondi-Sachs framework Geiller24 .
@@
The notion of mass in general relativity is a quite subtle issue. Iyer and Wald established an effective framework to naturally introduce conserved quantities, including the mass Iyer94 . This formalism works well in various diffeomorphism-covariant theories of gravity. Within the Iyer-Wald formalism, it was shown that the Arnowitt-Deser-Misner (ADM) mass is simply the Hamiltonian conjugate to the asymptotic time translation in spacetimes that are asymptotically flat at spatial infinity. To study the conserved quantities at null infinity, Wald et al. introduced a modified Hamiltonian and discussed the BMS charges in the conformal (unphysical) spacetime Wald2000 ; Grant2022 . In the current paper, we reexamine the Noether charge associated with the asymptotic time translation at null infinity in a direct way in the physical spacetime and no conformal compactification is required. It is shown that, for smooth asymptotically flat spacetimes, the Noether charge associated with the asymptotic time translation turns out to be the Bondi mass in the BS framework and the Newman-Unti (NU) mass respectively in the Newman-Penrose formalism. Moreover, the notion of mass at null infinity in polyhomogeneous spacetimes is discussed.
@@
Gravitational wave memory is a significant prediction of general relativity which was firstly found by Zel’dovich and Polnarev Zel74 . This type of memory is produced by the change of quadrupole moments and is referred to as the linear memory. In the 1990s, Christodoulou and Frauendiener found the nonlinear memory of gravitational waves Christo91 ; Frau92 . Recently, for smooth asymptotically flat spacetimes, Nichols et al. proposed a BMS method to calculate the gravitational wave memory Nichols17 ; Khera21 . Cao et al. applied the BMS method to calculate the gravitational memory produced by binary black hole coalescence Cao21-1 ; Cao21-2 . One of the essential steps in the BMS method is deriving the balance law (Cao21-1, , Eq.(15)), which relates the mass aspect to the asymptotic shear and thereby shows how gravitational radiation reduces the Bondi mass. The gravitational wave memory and the balance law for polyhomogeneous spacetimes are discussed in this paper. It is shown that the appearance of the logarithmic terms does not affect the balance law and it remains unchanged as the one that for smooth asymptotically flat spacetimes.
@@
We adopt the geometric units $c=G=1$ throughout this paper. Following the convention established in the Newman-Penrose formalism NP62 , we choose to use the Landau-Lifshitz timelike convention for the signature of the spacetime metric, (+,,,)(+,-,-,-). However, we shall make one exception to this convention in Section 4 to have the metric signature (,+,+,+)(-,+,+,+) which Bondi et al. used. This is because it is more convenient for readers to revisit the relevant formulae for smooth asymptotically flat spacetimes in Bondi62 ; Sachs62 ; CWYcmp2021 .
2 Asymptotic behaviour of Newman-Penrose quantities
@@
We follow the notions in NP62 . In addition to the null vector $l^{a}$, another null vector $n^{a}$ and the other two complex null vectors $m^{a},\bar{m}^{a}$ can be written as
@@
$\kappa=\epsilon=\pi=0,\ \rho=\bar{\rho},\ \tau=\bar{\alpha}+\beta,$ (2.3)
where $\kappa,\epsilon,\pi,\rho,\tau,\alpha$ and $\beta$ are spin coefficients NP62 . This gauge condition follows from the fact that $l_{a}=\nabla_{a}u$ is a gradient field and it implies that the tetrad vectors $m^{a}$ and $n^{a}$ are chosen to be parallel-propagated along $l^{a}$. Note that the gauge choice (2.3) we adopt here is slightly different with the one used in (Kroon99, , Eqs.(13) and (14)).
Within the Newman-Penrose formalism, the 38 NP equations (including the metric equations, the spin coefficient equations and the Bianchi identities) can be solved order-by-order. For this purpose, we need $\Psi_{0}$ on an initial null hypersurface $\mathcal{N}_{0}$ Kroon98 .
The most general polyhomogeneous form for the physically reasonable $\Psi_{0}$ Kroon99 is
@@
When solving the NP equations, one usually needs to calculate both the derivatives and the products of the polyhomogeneous functions Kroon98 . Let $h$ and $g$ be two polyhomogeneous functions as
@@
where $\rho_{i,k}$ and $\sigma_{i,k}$ are functions of $u,\theta$ and $\varphi$. We will denote by #\# the degree of a polynomial in $z$ and use the convention that the zero polynomial has degree -$\infty$.
In the following, we will solve the NP equations order-by-order. The procedure is similar to the one done in NP62 ; Kroon99 . We only give the details for the solving process of the first two equations in the NP hierarchy which are
@@
$\begin{split}\sigma=&\frac{\sigma_{2}(z)}{r^{2}}+\frac{\sigma_{3}(z)}{r^{3}}+\frac{\sigma_{4}(z)}{r^{4}}+\cdots,\\ &{}^{\#}\sigma_{2}=N_{3}+1,\ ^{\#}\sigma_{3}=N_{4},\ ^{\#}\sigma_{4}=\max\{3N_{3}+3,N_{5}\},\cdots,\\ &\sigma_{2}=\sigma_{2,0}(u,\theta,\varphi)+\sum_{i=1}^{N_{3}+1}\sigma_{2,i}(\theta,\varphi)z^{i},\\ &\sigma_{2,j}=\frac{1}{j}\Psi_{0}^{3,j},\ (j=1,2,\cdots,N_{3}+1).\end{split}$ (2.29)
$\begin{split}\rho=&-\frac{1}{r}+\frac{\rho_{3}(z)}{r^{3}}+\frac{\rho_{4}(z)}{r^{4}}+\cdots,\\ &{}^{\#}\rho_{3}=2N_{3}+2,\ ^{\#}\rho_{4}=N_{3}+N_{4}+1,\cdots,\\ &\rho_{3}=\sum_{j=0}^{2N_{3}+2}\rho_{3,j}z^{j},\\ &\rho_{3,2N_{3}+2}=-\sigma_{2,N_{3}+1}\bar{\sigma}_{2,N_{3}+1},\\ &\rho_{3,j}=(j+1)\rho_{3,j+1}-\sum_{k=0}^{j}\sigma_{2,k}\bar{\sigma}_{2,j-k},\ (j=2N_{3}+1,\cdots,1,0).\end{split}$ (2.30)
@@
$\begin{split}\alpha=&\frac{\alpha_{1}}{r}+\frac{\alpha_{2}(z)}{r^{2}}+\cdots,\ \ \ ^{\#}\alpha_{1}=0,\ ^{\#}\alpha_{2}=N_{3}+1,\cdots,\\ &\alpha_{1}=\alpha_{1,0}=-\frac{\cot\theta}{2\sqrt{2}},\\ &\alpha_{2}=\sum_{j=0}^{N_{3}+1}\alpha_{2,j}z^{j},\\ &\alpha_{2,N_{3}+1}=-\beta_{1,0}\bar{\sigma}_{2,N_{3}+1},\ \ \beta_{1,0}=\frac{\cot\theta}{2\sqrt{2}},\\ &\alpha_{2,j}=(j+1)\alpha_{2,j+1}-\beta_{1,0}\bar{\sigma}_{2,j},\ (j=N_{3},\cdots,1,0).\ \ \ \ \ \ \ \ \ \ \ \ \ \end{split}$ (2.31)
@@
$\begin{split}\beta=&\frac{\beta_{1}}{r}+\frac{\beta_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\beta_{1}=0,\ ^{\#}\beta_{2}=N_{3}+1,\cdots,\\ &\beta_{1}=\beta_{1,0}=\frac{\cot\theta}{2\sqrt{2}},\\ &\beta_{2}=\sum_{j=0}^{N_{3}+1}\beta_{2,j}z^{j},\\ &\beta_{2,N_{3}+1}=-\alpha_{1,0}\sigma_{2,N_{3}+1},\\ &\beta_{2,j}=(j+1)\beta_{2,j+1}-\alpha_{1,0}\sigma_{2,j}-\Psi_{1}^{3,j},\ (j=N_{3},\cdots,1,0).\ \ \ \end{split}$ (2.32)
$\begin{split}\tau=&\frac{\tau_{2}}{r^{2}}+\frac{\tau_{3}}{r^{3}}+\cdots,\ \ \ ^{\#}\tau_{2}=N_{3},\cdots,\\ &\tau_{2}=\sum_{j=0}^{N_{3}}\tau_{2,j}z^{j},\\ &\tau_{2,j}=\bar{\alpha}_{2,j}+\beta_{2,j},\ (j=0,1,\cdots,N_{3}).\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \end{split}$ (2.33)
$\begin{split}\lambda=&\frac{\lambda_{1}}{r}+\frac{\lambda_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\lambda_{1}=0,\ ^{\#}\lambda_{2}=N_{3}+1,\cdots,\\ &\lambda_{1}=\dot{\bar{\sigma}}_{2,0},\\ &\lambda_{2}=\sum_{j=0}^{N_{3}+1}\lambda_{2,j}z^{j},\\ &\lambda_{2,N_{3}+1}=-\mu_{1,0}\bar{\sigma}_{2,N_{3}+1}=\frac{1}{2}\bar{\sigma}_{2,N_{3}+1},\\ &\lambda_{2,j}=(j+1)\lambda_{2,j+1}+\frac{1}{2}\bar{\sigma}_{2,j},\ (j=N_{3},\cdots,1,0).\ \ \ \ \ \ \ \ \ \ \ \ \ \end{split}$ (2.34)
$\begin{split}\mu=&\frac{\mu_{1}}{r}+\frac{\mu_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\mu_{1}=0,\ ^{\#}\mu_{2}=N_{3}+1,\cdots,\\ &\mu_{1}=-\frac{1}{2},\\ &\mu_{2}=\sum_{j=0}^{N_{3}+1}\mu_{2,j}z^{j},\\ &\mu_{2,N_{3}+1}=-(\lambda_{1,0}\sigma_{2,N_{3}+1}+\Psi_{2}^{3,N_{3}+1}),\\ &\mu_{2,j}=(j+1)\mu_{2,j+1}-\sigma_{2,j}\lambda_{1,0}-\Psi_{2}^{3,j},\ (j=N_{3},\cdots,1,0).\end{split}$ (2.35)
@@
$\begin{split}\gamma=&\frac{\gamma_{2}}{r^{2}}+\frac{\gamma_{3}}{r^{3}}+\cdots,\ \ \ \ ^{\#}\gamma_{2}=N_{3}+1,\cdots,\\ &\gamma_{2}=\sum_{j=0}^{N_{3}+1}\gamma_{2,j}z^{j},\\ &\gamma_{2,N_{3}+1}=-\frac{1}{2}\Psi_{2}^{3,N_{3}+1},\\ &\gamma_{2,j}=\frac{1}{2}\bigg[(j+1)\gamma_{2,j+1}-\alpha_{1,0}\tau_{2,j}-\beta_{1,0}\bar{\tau}_{2,j}-\Psi_{2}^{3,j}\bigg],(j=N_{3},\cdots,1,0).\end{split}$ (2.36)
$\begin{split}\nu=&\frac{\nu_{1}}{r}+\frac{\nu_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\nu_{1}=0,\ ^{\#}\nu_{2}=N_{3}+1,\cdots,\\ &\nu_{1}=-\Psi_{3}^{2,0}=\eth\dot{\bar{\sigma}}_{2,0},\\ &\nu_{2}=\sum_{j=0}^{N_{3}+1}\nu_{2,j}z^{j},\\ &\nu_{2,N_{3}+1}=-\frac{1}{2}\Psi_{3}^{3,N_{3}+1},\\ &\nu_{2,j}=\frac{1}{2}\bigg[(j+1)\nu_{2,j+1}+\frac{1}{2}\bar{\tau}_{2,j}-\lambda_{1,0}\tau_{2,j}-\Psi_{3}^{3,j}\bigg],\ (j=N_{3},\cdots,1,0).\end{split}$ (2.37)